How do you solve (x-9)(x+1)(x-1)=0
Found 2 solutions by stanbon, Angel73:
Answer by stanbon(75887)

solve (x-9)(x+1)(x-1)=0
----------------
x = 9 or x = -1, or x = 1
===============================
Cheers,
Stan H.

Answer by Angel73(1)

you set each one equal to zero so u the equations would be (x-9)=0 (x+1)=0 (x-1)=0 then u solve each one as a seperate equation and u get X is (9,-1&1) in these equations u usually get more than 1 answer.